# ,, thanks: These two authors contributed equally to this work thanks: These two authors contributed equally to this work

### v1

@@ Abstract
where $x_{k,a}$ and $n_{k,i}$ are the orbital and isospin coordinates for the $k$th particle, respectively. The power of the spinor $\psi$ of each particle appearing in Eq.(17) is equal to l=m(N1)l=m(N-1), hence, the filling should be $\nu=N/d(l,0)$ with $d(l,0)$ defined in Eq.(3) approaching $1/(m^{3}N^{2})\to 0$ as $N\to\infty$. The parent Hamiltonian to this wavefunction is constrained in channel $\bigoplus\limits_{n_{1}+n_{2}\geq l-m+1}(n_{1}+n_{2},n_{1}-n_{2})_{so5}$. There is no well-defined thermodynamic limit for the Jastrow-type wavefunction.
In conclusion, we have studied the 4D incompressible quantum liquid based on the Landau level under an SU(2) monopole gauge potential, specifically using the determinant-Laughlin and Jastrow wavefunctions. The exact Hamiltonian for these wavefunctions was constructed within the pseudo-potential formalism. The numerical calculations show that the unique ground state is an SO(5) singlet. In addition, the ground states with a quasi-hole remain zero energy. In comparison, the ground states with a quasi-particle exhibit a finite gap. The liquid-like behavior is further supported by the pairing distribution function. However, direct numerical analysis of both determinant-Laughlin and Jastrow wavefunctions for larger particle numbers turn out to be prohibitive due to the exponentially growing dimensions of the Hilbert space. Nevertheless, the recent experimental realization of the few-electron Coulomb liquid Shaju et al. (2025) also provides a motivation for studying few-particle strongly correlated systems. Another intriguing avenue is to explore the statistical properties of excitations in the 4D FQHE, particularly investigating loop or membrane statistics in higher dimensions Congjun Wu and Zhang (2010); Wang and Levin (2014); Chen et al. (2023b); Bernevig et al. (2002); Wang et al. (2019); Huang et al. (2024); Huxford et al. (2024). Furthermore, studying the entanglement spectrum Kitaev and Preskill (2006); Levin and Wen (2006); Li and Haldane (2008) would offer valuable insights into these states and warrants further investigation.

### v2

@@ Abstract
The power of the spinor $\psi$ of each particle appearing in Eq.(17) is equal to l=m(N1)l=m(N-1), hence, the filling should be $\nu=N/d(l,0)$ with $d(l,0)$ defined in Eq.(3) approaching $1/(m^{3}N^{2})\to 0$ as $N\to\infty$. The parent Hamiltonian to this wavefunction is constrained in channel $\bigoplus\limits_{n_{1}+n_{2}\geq l-m+1}(n_{1}+n_{2},n_{1}-n_{2})_{so5}$. There is no well-defined thermodynamic limit for the Jastrow-type wavefunction.
In conclusion, we have studied the 4D incompressible quantum liquid based on the Landau level under an SU(2) monopole gauge potential, specifically using the determinant-Laughlin and Jastrow wavefunctions. The parent Hamiltonian for these wavefunctions was constructed within the pseudo-potential formalism. The numerical calculations show that the unique ground state is an SO(5) singlet. In addition, quasi-hole states remain at zero energy. In contrast, quasi-particle excitations are separated by a finite gap. The liquid-like behavior is further supported by the pair distribution function. However, direct numerical analysis of both determinant-type Laughlin and Jastrow wavefunctions for larger particle numbers is prohibitive due to the enormous size of the Hilbert space. For $N=10$, the next accessible system size, the Hilbert space has a dimension of about $10^{11}$.Nonetheless, the recent experimental demonstration of the few-electron Coulomb liquid Mukherjee et al. (2022); Lunt et al. (2024); Shaju et al. (2025) provides a promising platform for the experimental validation of our theoretical framework. Another intriguing avenue is to explore the statistical properties of excitations in the 4D FQHE, particularly investigating loop or membrane statistics in higher dimensions Congjun Wu and Zhang (2010); Wang and Levin (2014); Chen et al. (2023b); Bernevig et al. (2002); Wang et al. (2019); Huang et al. (2024); Huxford et al. (2024); Agarwal et al. (2025). Furthermore, studying the entanglement spectrum Kitaev and Preskill (2006); Levin and Wen (2006); Li and Haldane (2008) would offer valuable insights into these states and warrants further investigation.